OpenStudy (anonymous):

solve using elimination y=3x+4 y=2x+2

OpenStudy (anonymous):

x=-2/3 y=2/3

OpenStudy (anonymous):

Need work?

OpenStudy (anonymous):

yes please:)

OpenStudy (anonymous):

they are both y, set the right hands sides equal

OpenStudy (anonymous):

3x+4 -2(2x+2) eliminate the 4 ------- 3x+4 -4x-4 --------- y=-x then right solve it from here right?

OpenStudy (anonymous):

i got -2 idk if its right

OpenStudy (anonymous):

its almost like a regular subtraction problem but you have to make 1 term (any 1 you want) equal to the additive inverse to eliminate it 3x+4 -2(2x+2) eliminate the 4 i find is the easiest -------

OpenStudy (anonymous):

is it -2?

OpenStudy (anonymous):

wow that is so much work \[y=3x+4=2x+2\] subtract \(2x\) from both sides get \[x+4=2\] subtract 4 from both sides get \[x=-2\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thanks All =)

OpenStudy (anonymous):

there were other ways but it was elimination

OpenStudy (anonymous):

if you see \[y=\text{expression}\] and \[y=\text{another expression}\] it is correct to set them equal and solve, and frankly somewhat silly to use another method. (i am not saying you weren't asked to, just that it is silly)